OpenStudy (anonymous):

convert each degree measure to radian measure. A: -200 degrees b: 455 degrees C:-30 degrees

OpenStudy (solomonzelman):

pi=180

OpenStudy (ray10):

take each value in degrees, and multiply it by \[\large \frac{ \pi }{ 180 }\]

OpenStudy (ray10):

\[\large e.g. >30*\frac{ \pi }{ 180 } = \frac{ \pi }{ 6 }\]

OpenStudy (anonymous):

thanks!

OpenStudy (ray10):

you're welcome :) I hoped it helped!

OpenStudy (anonymous):

It did :D

OpenStudy (ray10):

great to hear! All the best in future conversion questions. I'll help if I can :)